# The Nonlinear Filter Model of Stream Cipher Redivivus

### v6

@@ 6 Conclusion
We described a construction of balanced Boolean functions which has several provable properties, namely very high nonlinearity, acceptable algebraic resistance, and is efficient to implement in hardware. Using such Boolean functions, we proposed concrete constructions of the nonlinear filter model for stream ciphers targeted at different security levels. Gate count estimates for the stream cipher proposals show that the circuit sizes compare well with famous stream ciphers at the 80-bit and the 128-bit security levels, while for higher security levels, we do not know of any stream cipher with lower gate count estimates.
@@ Acknowledgement
We thank Subhadeep Banik, Tim Beyne, Pierrick Méaux, Will Meier, Michiel Verbauwhede, and Bin Zhang for their comments on earlier versions of the paper. Deng Tang provided us with a program written by Simon Fischer which we have used for computing fast algebraic immunity. We thank both of them.

### v7

@@ 6 Conclusion
We described a construction of balanced Boolean functions with several provable properties, namely very high nonlinearity, acceptable algebraic resistance, and enabling efficient hardware implementation. Using such Boolean functions, we proposed concrete constructions of the nonlinear filter model for stream ciphers targeted at different security levels. Gate count estimates for the stream cipher proposals show that the circuit sizes compare well with famous stream ciphers at the 80-bit and the 128-bit security levels, while for higher security levels, we do not know of any stream cipher with lower gate count estimates.
@@ Acknowledgement
We thank Subhadeep Banik, Tim Beyne, Pierrick Méaux, Willi Meier, Michiel Verbauwhede, and Bin Zhang for their comments on earlier versions of the paper. We also thank the anonymous reviewers for helpful comments. Deng Tang provided us with a program written by Simon Fischer which we have used for computing fast algebraic immunity. We thank both of them.